# Writing – original draft, Data curation, Formal analysis, Methodology, Conceptualization

### v1

@@ Abstract
The COVID-19 pandemic underscored the critical need for rapid epidemic trend identification and effective intervention strategies to mitigate disease progression and its socio-economic impact. Concurrent with emerging threats, endemic diseases like dengue continue to strain healthcare systems, particularly in populous, economically challenged nations. This paper introduces AVSim (Airborne and Vectorborne Simulator), an agent-based model designed to provide granular insights for optimizing resource allocation within existing healthcare management frameworks. AVSim leverages realistic human mobility and behavioral patterns to simulate disease propagation within a detailed, scalable environment encompassing homes, schools, hospitals, and commercial venues. Human movement is modeled based on occupational and behavioral patterns, including age-specific activities. The simulator incorporates age- and environment-specific disease outcomes, host-host and host-vector interactions, and multiple disease stages, including mild, severe, and critical phases. Immunity, quarantine, and hospitalization are also modeled. Furthermore, AVSim supports tracing the path of disease spread, providing micro-level insights into transmission dynamics. Implemented in Python, AVSim offers flexibility and extensibility, enabling users to create highly customized scenarios for airborne and vector-borne disease modeling. Case studies demonstrating AVSim’s application to COVID-19 and dengue illustrate its potential for generating actionable epidemic insights, thereby enhancing public health planning and response.
keywords:
[…]
{highlights}
Human mobility modeling using machine learning and probabilistic models.
Unveiling hidden behavioral patterns via unsupervised clustering and graph theory.
Agent-based model with realistic scalable environments and transportation systems.
Simulates airborne and vector-borne disease spread dynamics.
Provides micro-level insights for public health planning and response.
1 Introduction
@@
The COVID-19 pandemic lasted over three years, resulting in more than 700 million cases and claiming over 7 million lives [1]. This crisis highlighted the critical importance of timely interventions and coordinated efforts by individuals, healthcare systems, and governments to control disease outbreaks and prevent them from escalating into pandemics. Beyond COVID-19, vector-borne diseases, transmitted by living organisms that carry infectious pathogens between humans or from animals to humans, pose a significant public health challenge, causing over 700,000 deaths annually. For instance, In 2024, the World Health Organization (WHO) estimated approximately 249 million global cases of malaria and 100–400 million cases of dengue each year [2]. Many vector-borne and airborne diseases, including those similar to COVID-19, are preventable through protective measures and effective community mobilization. These preventable diseases underscore the need for continued vigilance, innovative interventions, and global collaboration to mitigate the risk of future outbreaks and safeguard public health.
Numerous methods have been employed to model disease spread and uncover patterns to better understand disease propagation so that effective disease control strategies could be adopted efficiently while minimizing the net burden. These models can be broadly classified into two categories: compartment models and agent-based models (ABMs). Compartment models are relatively simple and computationally efficient. Within the compartmental modeling framework, the Susceptible-Exposed-Infectious-Recovered (SEIR) model has been widely used to study epidemics and pandemics [3]. These models classify individuals by their disease status and use differential equations to describe transitions between these states. Although they can give an overview of the disease’s spread, they fail to capture many important details. Despite this, numerous attempts exist to model disease spread with such compartment models. Shaobo He et al. [4] used an SEIR compartment model based on the data from Hubei province (China) to show the evolution of the COVID-19 epidemic. Jonas Dehning et al. [5] focused on COVID-19 spread in Germany and used a SIR model to quantify the effect of interventions. Giulia Giordano et al. [6] proposed a new extended compartment model, SIDARTHE (Susceptible-Infected-Diagnosed-Ailing-Recognized-Threatened-Healed-Extinct), that predicts the course of the epidemic to help plan an effective control strategy.
Agent-based models, on the other hand, are more complex and able to deliver micro-level details about each agent in the simulation. They have the capability to model interactions between agents or entities in a defined environment. Each agent follows a set of well-defined rules, allowing ABMs to capture emergent behaviors that arise from these interactions. This makes them particularly useful for modeling complex systems in fields like epidemiology [7, 8], urban planning [9], ecology [10], and molecular-level biological [11, 12] systems. The ability to incorporate fine-grained details enhances decision-making, especially for policymakers seeking accurate predictions.
The use of ABMs in epidemiology has grown significantly during the COVID-19 pandemic. Models like Covasim [7], PDSIM [13], and OpenABM [8] have been widely used to analyze disease dynamics and inform public health responses. Some benchmark ABMs are mentioned and compared with our proposed ABM in Table 1.
Additionally, there has been increasing research on the dynamic analysis of vector-borne diseases. One notable approach by Carrie Manore et al. [14] employs an ABM to analyze the propagation of mosquito-borne diseases. This study introduces the network-patch method, utilizing differential equations to model mosquito density. Similarly, Imran Mahmood et al. [15] use an ABM to compute vector density based on the reproductive behavior of vectors. However, their model does not incorporate realistic movement patterns, limiting its ability to provide real-world insights. Another study applied a fuzzy modeling approach to develop an epidemic model for mosquito-borne diseases [16].
A comprehensive modeling and analysis tool, such as ABM or otherwise, helps to be better prepared for an epidemic situation. For example, during the initial stages of the COVID pandemic, there were many unknowns and uncertainties resulting in confusion, stressing of healthcare resources, as well as burdening the public excessively due to the control strategies. Many of the initial responses involved in reactive measures such as contact tracing and mass-scale quarantining and lock-downs. A comprehensive computational model would, on the other hand, complement the existing diverse healthcare systems and practices of countries and regions by making it possible to "play out" different scenarios using the best of data and knowledge available and/or learned assumptions on what may happen under different cases. Armed with all these possible outcomes, healthcare resources could be targeted effectively and efficiently so that maximum outcomes could be attained.
@@
Effective intervention and prevention strategies in epidemiology require modeling disease spread under realistic conditions. Human mobility, behavioral patterns, and public modes of transportation significantly influence the speed and extent of pathogen transmission [13]. Different behaviors and transport modes affect disease spread in varied ways. However, many models either overlook these variations [17, 18] or rely on oversimplified assumptions, such as treating transportation as random interactions aggregated with activities like shopping or social events [7, 19], failing to reflect real-world complexities.
To address the limitations of existing methods, we propose an agent-based model for vector-borne and airborne diseases, named AVSim, that uses real-world data and accounts for real-world conditions. Our framework learns and models different human mobility and behavioral patterns, simulating agents based on occupation and behavior at a granular level. It incorporates a realistic environment and public transportation model to simulate disease spread. The proposed model is customizable and extendable, offering a flexible framework adaptable to various diseases and transmission modes, thus enhancing its applicability to emerging public health challenges. A preliminary version of this work was presented in [20, 21]. In this article, we expand on our previous findings by introducing additional experiments, deeper analysis, and improved methodologies. We further demonstrate the framework’s use in simulating COVID-19 and dengue propagation through an extended SEIR model within an ABM.
@@
Table 1: Comparison of disease simulators found in the literature with ours.
@@
Use of real mobility data
@@
Tracing Path of Disease Spread
@@
ABM for vaccine efficacy’s influence on public health projections [23]
@@
AVSim (Our Approach)
@@
People meet people [27]
COVID-town [28]
An Agent-based modeling of COVID-19 [29]
Social bubble vanpooling (SBV) [30]
AVSim (Our Approach)
2 Material and methods
@@
The proposed AVSim consisted of two major stages. The first stage was the data generation stage, where human mobility and behavioral patterns were identified and generated using real-world data. This sets the foundation for the second stage of AVSim: the simulation of the ABM. As shown in Fig. 1, the ABM simulation stage comprised four main components: the environment, vector patch, transportation, and agent generation. However, for simulating vector-borne diseases, only the environment, vector patch, and agent generation components were used, as transportation had no significant impact on disease spread. Conversely, for airborne diseases, only the environment, transportation, and agent generation components were utilized. Each of these stages will be explained in detail in the following sections.
@@
2.1 Modeling human behaviour
@@
To accurately model disease spread in a human population, it was essential to replicate realistic human motion patterns within the simulation environment. In the context of human motion, various factors affected:
  • Time-dependent location visits: Individuals visit specific locations based on the time of day.
  • Duration of stay: The length of time spent at a location varies depending on the time of day and the nature of the location.
@@
  • Profession-dependent behavior: The types of locations visited and the associated durations differ based on individuals’ professions.
To address these factors, In the first stage of AVSim, we proposed a method for generating realistic human motion patterns using real-world GPS data. This approach enabled us to delve deeper into human mobility and behavioral dynamics, uncovering insights that were not immediately apparent. The following subsections outline the data acquisition process and the methods used for pattern recognition.
2.1.1 Tracking human movement and data acquisition
The first step in modeling human motion was gathering a large and diverse dataset of real-world movement patterns. To achieve this, GPS tracking of human agents was conducted using a smartphone application. The dataset was sourced from 100 voluntary participants residing in the Kandy district of Sri Lanka, each representing one of 13 distinct professions. The participants’ GPS coordinates, altitude, and timestamps were recorded at 5-minute intervals over 21 days. This duration was chosen as the optimal time frame for capturing representative mobility patterns while remaining manageable for participants. A longer duration could have led to participant fatigue or reduced engagement, affecting data quality. To enhance diversity and fairness, participants within a single profession were carefully selected to reflect varying factors such as age, gender, and the relative distance between their workplace and residence.
Figure 2: An example of accepted and unaccepted patterns based on daily data availability.
Before enrollment, all participants were well informed about the scope of the data being collected and its intended use. Informed verbal and written consent were obtained from each participant, with students requiring additional consent from a parent or guardian and their educational institution. Upon consent, participants were provided with the smartphone application and instructed to keep their smartphones active and within reach throughout the study. Data privacy and confidentiality were maintained at all stages, and all data was anonymized.
@@ 2.1.1 Tracking human movement and data acquisition
Ethical approval: The ethical clearance for this study was obtained from the Ethical Review Committee, Faculty of Arts, University of Peradeniya, Sri Lanka, (ARTS/ERC/ 2021/01 and 18 September 2021) with the support of the Department of Sociology. Administrative clearance was obtained by the Ministry of Home Affairs of Sri Lanka, relevant Divisional Secretariate (DS) offices, and relevant Grama Niladari (GN) offices.
To ensure data quality, daily datasets were reviewed for completeness, and 21 complete daily datasets were required per participant. A day was considered complete if at least 80% of the expected 288 data points (recorded every 5 minutes) were present and evenly distributed throughout the day, minimizing gaps in coverage. Days with more than 20% missing data (57 data points) or significant periods of consecutive data loss were excluded. As a result, data collection for a participant typically extended beyond 21 days to ensure 21 days of complete data. Fig. 2 shows an example of accepted and discarded data based on its availability throughout the day.
2.1.2 Human mobility patterns identification
The GPS data was recorded in the time and coordinate domain (longitude and latitude), making it specific to the geographical environment where it was collected, in this case, Kandy, Sri Lanka. However, human motion patterns are generally consistent across different geographical locations. For example, a teacher’s typical working day will not depend on the geographical location. They will be at the resting place in the morning, leave for work, maybe visit some places in the evening, and return to the resting place at the end of the day. To make the data applicable beyond the specific geographic context, we transformed it into a time and location domain that removes the dependency on physical coordinates. This transformation generalizes the data, enabling broader applicability and analysis of human motion patterns.
To identify the locations visited by each person, the GPS data for all participants over the 21 days was plotted on a map of Sri Lanka. Fig. 3 illustrates the raw GPS data when mapped geographically. The data reflects the places visited and the movements during transportation, such as commuting between locations.
@@ 2.1.2 Human mobility patterns identification
When zooming in on these denser clusters, which occur at locations where the person pauses, additional sub-clusters emerge within them, revealing further granularity.
@@
Fig. 3 demonstrates this inherent pattern in human motion. This pattern highlights a hierarchical structure in the locations visited, with primary locations further divided into more specific destinations.
For this study, our primary interest lies in the denser clusters of GPS data, which indicate areas where participants spent more time, such as workplaces, homes, and other specific locations.
To identify these clusters, a clustering method needs to be deployed. In this study, we employed the Density-Based Spatial Clustering of Applications with Noise (DBSCAN), which produced effective results. DBSCAN is a popular density-based clustering algorithm known for detecting clusters of varying densities while being robust to noise. References  [31] and  [32] proved to be useful in implementing the DBSCAN algorithm. Unlike K-Means clustering and Gaussian Mixture Models, DBSCAN does not require the number of clusters to be specified in advance. Instead, it determines the number of clusters based on the data, which makes it well-suited for identifying clusters in GPS data. DBSCAN classifies points into three categories: core points, border points, and noise points. Core and border points form the clusters, while noise points are discarded. As a result, transportation-related data is treated as noise and removed from the analysis. The DBSCAN algorithm requires two initial parameters:
  1. 1.
    epsilon ($\epsilon$) defines the maximum distance between two points in a cluster.
  2. 2.
    minPts, the minimum number of points required for a point to be considered a core point of a cluster.
To determine the most suitable values for these parameters, we experimented with different values of $\epsilon$ and minPts. Through an iterative process of testing and evaluating clustering performance, we found that setting $\epsilon$ = 5 and minPts = 10 produced the most meaningful and consistent clusters. These values effectively distinguished visited locations while minimizing noise, making them optimal for our specific application.
Since the primary goal was to convert time-coordinates domain data into the time-location domain, the next step involved identifying the numbered clusters as real-world locations. For this, commonly visited locations such as "Home," "School," and "Bank" were made into a list and grouped into broader categories like "Residential Zone," "Education Zone," and "Commercial and Financial Zone". Each cluster was then labeled based on these predefined locations and zones.
Figure 4: Clustering GPS data using DBSCAN algorithm.
For example, if a cluster represented an individual’s home, it was labeled as "Home." Similarly, clusters corresponding to areas around the individual’s neighborhood were labeled as "Residential Zone."
During this process, some clusters could not be mapped to identifiable real-world locations. This may have occurred for several reasons, such as GPS data drift, a phenomenon where the reported GPS position shifts over time despite the device being stationary, or network connectivity issues that affected the accuracy of location tracking. Additionally, even if the individual visited the area corresponding to a cluster, it might not match any meaningful location in the predefined list. These clusters were treated as outliers and excluded from the analysis.
The original dataset was recorded at a 5-minute resolution with some missing data points. However, the emulator (described in subsequent sections) operates at a 1-minute resolution. To bridge this gap, the data was reconstructed using Zero Order Hold (ZOH) reconstruction. This method was deemed sufficient because of several key factors. First, it was reasonable to assume that a person’s location remained approximately constant within the 5-minute intervals between recorded points. Additionally, the dataset had been carefully preprocessed to ensure that missing data was minimal, allowing interpolation to address these gaps with negligible inaccuracies.
The reconstructed dataset with a 1-minute resolution is visualized in Fig.  5. It is important to note that transportation-related data points were excluded during the DBSCAN clustering process and are, therefore, not present in the final dataset. This dataset was used to model human movements in the proposed AVSim.
2.1.3 Human behavioural patterns identification
In this study, GPS data was collected from individuals belonging to various professional classes. This data was used to uncover the hidden structure of society based on occupation. Within each professional class, sub-professional class behaviors can be identified, as shown in Fig.  1.
@@ 2.1.3 Human behavioural patterns identification
The Spectral Clustering method, which utilizes spectral graph theory [33], was employed to uncover hidden behavior patterns or sub-classes within each occupational class. Spectral clustering is an algorithm that considers the global structure of a dataset and determines the optimal number of clusters within it.
Although there is a lot of literature on spectral clustering within machine learning, the sources  [34] and  [35] proved to be of great value when applying spectral clustering to unlock hidden patterns in the dataset.
For the clustering process, locations needed to be encoded numerically. A 6-digit binary encoding was used, where the first three digits uniquely represent the zones, and the last three digits uniquely identify the locations within those zones, as defined in Section 2.1.2. This encoding ensured a clear separation between distinct locations, while locations within the same zone had less separation, reflecting their relative proximity or similarity. This compact numerical representation of locations facilitated effective clustering and produced acceptable results, as elaborated in Section 3.1.
Let us consider a day as a data point. Then we can represent our data set $(X)$ as, $X=x_{1},x_{2},x_{3},\dots,x_{n}$ and the notion of similarity between $x_{i}$ and $x_{j}$ as $S_{ij}>0$.
To define similarity, we have chosen a Gaussian kernel, as shown below.

### v2

# AVSim - Realistic Simulation Framework for Airborne and Vector-Borne Disease Dynamics

@@ Abstract
Computational disease modeling plays a crucial role in understanding and controlling the transmission of infectious diseases. While agent-based models (ABMs) provide detailed insights into individual dynamics, accurately replicating human motion remains challenging due to its complex, multi factorial nature. Most existing frameworks fail to model realistic human motion, leading to oversimplified and less realistic behavior modeling. Furthermore, many current models rely on synthetic assumptions and fail to account for realistic environmental structures, transportation systems, and behavioral heterogeneity across occupation groups. To address these limitations, we introduce AVSim, an agent-based simulation framework designed to model airborne and vector-borne disease dynamics under realistic conditions. A distinguishing feature of AVSim is its ability to accurately model the dual nature of human mobility (both the destinations individuals visit and the duration of their stay) by utilizing GPS traces from real-world participants, characterized by occupation. This enables a significantly more granular and realistic representation of human movement compared to existing approaches. Furthermore, spectral clustering combined with graph-theoretic analysis is used to uncover latent behavioral patterns within occupations, enabling fine-grained modeling of agent behavior. We validate the synthetic human mobility patterns against ground-truth GPS data and demonstrate AVSim’s capabilities via simulations of COVID-19 and dengue. The results highlight AVSim’s capacity to trace infection pathways, identify high-risk zones, and evaluate interventions such as vaccination, quarantine, and vector control with occupational and geographic specificity.
Index Terms:
Agent-based models, social systems, epidemiology, unsupervised clustering, COVID-19, Dengue
I Introduction
@@
Disease modeling is crucial in modern society for predicting and managing infectious disease spread, enabling proactive public health policies. It optimizes resource allocation, informs targeted interventions, and enhances forethought for emerging epidemic threats. The need for robust computational frameworks to model epidemic dynamics is highlighted by the rapid spread of airborne and vector-borne diseases. Starting from December 2019, the COVID-19 pandemic lasted over three years, resulting in more than 700 million cases and claiming over 7 million lives [1]. This crisis highlighted the critical importance of timely interventions and coordinated efforts by individuals, healthcare systems, and governments to control disease outbreaks and prevent them from escalating into pandemics. Beyond COVID-19, vector-borne diseases, transmitted by living organisms that carry infectious pathogens between humans or from animals to humans, pose a significant public health challenge, causing over 700,000 deaths annually. For instance, in 2024, the World Health Organization (WHO) estimated approximately 249 million global cases of malaria and 100–400 million cases of dengue each year [2]. Many vector-borne and airborne diseases, including those similar to COVID-19, are preventable through protective measures and effective community mobilization. Modeling these disease dynamics helps to identify hot spots, optimize interventions and resource allocation, and further enable policymakers to implement proactive, data-driven strategies.
@@
Effective intervention and prevention strategies in epidemiology require modeling disease spread under realistic conditions. Human mobility, behavioral patterns, and public modes of transportation significantly influence the speed and extent of pathogen transmission [3]. Different behaviors and transport modes affect disease spread in varied ways. However, many models either overlook these variations [4, 5] or rely on oversimplified assumptions, such as treating transportation as random interactions aggregated with activities like shopping or social events [6, 7], failing to reflect real-world complexities. Furthermore, previous methods fail to capture realistic human motion patterns and underlying social dynamics. It is well known that modeling human behavior is a complex task due to the inherent complexity and unpredictability of human systems. Humans exhibit diverse behaviors influenced by factors like profession, age, culture, environment, and personal preferences. Not only that, but social dynamics arise from complex, nonlinear interactions among individuals and environments. Therefore, it is challenging to model this type of dynamical system and predict outcomes without a detailed understanding of the feedback loops and control mechanisms within the system.
@@
TABLE I: Comparison of disease simulators found in the literature with ours.
@@
Real Mobility Data Usage
@@
Disease Spread Tracing
@@
ABM for vaccine efficacy’s influence on public health [11]
@@
AVSim - Vector-borne model (Our Approach)
@@
Agent-based transmission model of COVID-19 [14]
AbCSim [15]
AVSim - Airborne model (Our Approach)
To overcome limitations in existing models, we present AVSim, an agent-based framework for simulating both airborne and vector-borne disease dynamics under realistic conditions. Unlike approaches based on synthetic mobility, AVSim derives agent behavior from real-world GPS data by identifying stay regions and abstracting them into reusable location types. It captures both spatial and temporal aspects of mobility, where individuals go and how long they stay, stratified by profession and context. Behavioral subtypes within occupations are uncovered via unsupervised spectral clustering, enabling nuanced agent modeling. These insights are embedded in a scalable simulation environment with hierarchical geography and integrated public transport. AVSim supports detailed disease modeling through an extended SEIR (Susceptible-Exposed-Infectious-Recovered) framework and is adaptable to a range of transmission modes. A preliminary version of this work appeared in [16, 17]; this article expands on it with improved methodologies, additional experiments, and validation through simulations of COVID-19 and dengue propagation.
In summary, we make the following contributions to the agent-based modeling community:
  • A novel human motion modeling approach that leverages GPS data and density-based spatial clustering to generate high-resolution daily agent trajectories.
  • Uncover heterogeneous behavioral patterns within the same occupation groups using unsupervised clustering and graph-theoretic analysis.
  • Introduction of an agent-based model with realistic, scalable environments and transportation systems.
  • A case study demonstrating the capability of the proposed ABM in simulating the spread of COVID-19 and dengue.
The rest of this paper is organized as follows: Section II reviews related work on disease modeling and human motion modeling using ABMs. Section III details the methods used to model and simulate human behavior within ABM. Section IV presents the results, discussion, and validation of our approach. Finally, Section V concludes by highlighting the significance of the proposed framework.
II Related Work
Numerous methods have been employed to model disease spread and uncover patterns to better understand disease propagation so that effective disease control strategies can be adopted. These models can be broadly classified into two categories: compartment models and agent-based models (ABMs). Compartment models are relatively simple and computationally efficient. Within the compartmental modeling framework, the SEIR model and its modified versions have been widely used to study epidemics and pandemics [18, 19, 20]. However, they fail to present granular, micro-level details about the simulation. ABMs, on the other hand, are more complex and able to deliver micro-level details about each agent in the simulation. They have the capability to model interactions between agents or entities in a defined environment. Each agent follows a set of well-defined rules, allowing ABMs to capture emergent behaviors that arise from these interactions. This makes them particularly useful for modeling complex systems in fields like epidemiology, healthcare systems, disaster response, manufacturing and automation, transportation and mobility, economic and market behavior  [6, 7, 21, 22, 23, 24]. The ability to incorporate fine-grained details enhances decision-making, especially for policymakers seeking accurate predictions. Some benchmark ABMs are mentioned and compared with our proposed ABM in Table I.
Additionally, there has been increasing research on the dynamic analysis of vector-borne diseases. One notable approach by Carrie Manore et al. [8] employs an ABM to analyze the propagation of mosquito-borne diseases. This study introduces the network-patch method, utilizing differential equations to model mosquito density. Similarly, Imran Mahmood et al. [9] use an ABM to compute vector density based on the reproductive behavior of vectors. However, their model does not incorporate realistic movement patterns, limiting its ability to provide real-world insights. Another study applied a fuzzy modeling approach to develop an epidemic model for mosquito-borne diseases [25]. Many of these models have modeled agents by representing them as nodes in a network  [6, 7]. In these networks, human behavior and interactions were modeled using probability distributions [26]. Although reasonably accurate, these methods fail to capture the full extent of human motion dynamics.
III Material and methods
@@
The proposed design of AVSim consisted of two major stages. The first stage was the data processing stage, where human mobility and behavioral patterns were identified and generated using real-world data. This sets the foundation for the second stage of AVSim: the design of the agent-based modeling system. As shown in Fig. 1, the ABM is comprised of four main components: the environment, vector patches, transportation network, and agents. Each of these components will be explained in detail in the following sections.
@@
III-A Modeling human behaviour
@@
To accurately model disease spread in a human population, it was essential to replicate realistic human motion patterns within the simulation environment. We identified the following factors that affect human motion:
  • Individuals visit specific locations based on the time of day.
  • The length of time spent at a location varies depending on the time of day and the nature of the location.
@@
  • The types of locations visited and the associated durations differ based on individuals’ professions.
To incorporate these factors, we introduced a novel GPS-based approach using real-world mobility data from 100 volunteers in Kandy, Sri Lanka, representing 13 professions. These profession classes were selected such that they capture the broader population of Kandy District. Location coordinates, altitude, and timestamps were recorded at 5-minute intervals over 21 days. Therefore, the dataset consisted of a total of 2100 days. Since the ABM simulator (described in subsequent sections) operates at a 1-minute resolution, the data was upsampled using Zero-Order Hold reconstruction.
@@
Ethical approval: This study was approved by the Ethical Review Committee, Faculty of Arts, University of Peradeniya, Sri Lanka (ARTS/ERC/2021/01, 18 September 2021), with support from the Department of Sociology. Administrative clearance was granted by Sri Lanka’s Ministry of Home Affairs, relevant Divisional Secretariat offices, and village officer divisions.
III-A1 Mobility pattern clustering
To generalize the geographically specific GPS data collected in Kandy, Sri Lanka, we transformed the coordinate-based (longitude/latitude) data into a time-location domain representation. This abstraction preserves essential mobility patterns while removing dependence on physical coordinates.
@@ III-A1 Mobility pattern clustering
To identify the locations visited by each person, the GPS data for all participants over the 21 days was plotted on a map of Sri Lanka. When zooming in on these denser clusters (which occur at locations where the person pauses), additional sub-clusters emerge within them, revealing further granularity. Fig. 2 illustrates this inherent pattern in human motion, highlighting a hierarchical structure in the locations visited, with primary locations further divided into more specific destinations.
@@
This study focuses on denser clusters of GPS data, indicating locations where participants spent more time. We used the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm to detect these clusters, as it handles varying densities and is robust to noise [27, 28].
DBSCAN requires two parameters: epsilon ($\epsilon$), the maximum distance between points (in meters), and minPts, the minimum number of points to form a cluster. We set ($\epsilon$) = 5 m and minPts = 10, which produced the most consistent and meaningful clusters.
To map clusters to real-world locations, we labeled them using predefined locations and zones. For example, if a cluster represented an individual’s home, it was labeled as ”Home.” Similarly, clusters corresponding to areas around the individual’s neighborhood were labeled as ”Residential Zone.”
Transportation points were excluded from clustering and do not appear in the final dataset. A part of the reconstructed dataset with a visual summary is shown in Fig. 3. The statistics needed to model human motion were derived from this dataset.
@@
III-A2 Behavioral patterns clustering
The processed dataset revealed occupation-based societal structures, with distinct behavioral sub-classes emerging within each profession, as shown in Fig. 1 (b). We applied Spectral Clustering [29] to identify these hidden patterns, building on methods from [30, 31].
For effective clustering, we numerically encoded locations using a 6-digit binary system where the first three digits uniquely identified zones and the remaining three digits specified locations within those zones. This encoding ensured a clear separation between distinct locations, while locations within the same zone had less separation, reflecting their relative proximity or similarity.
[…]
To define similarity, we have chosen a Gaussian kernel, as shown below. The Gaussian kernel assigns high similarity to close-by points, allowing them to cluster together, while maintaining clear separation between distinct groups.